OpenStudy (anonymous):

How do you find the second derivative of y=x^2-(1/x)+3?

OpenStudy (anonymous):

y' = 2x + x^-2 y'' = 2-2x^-3

OpenStudy (anonymous):

\[y=x^2-x^{-1}+3\] \[\delta y/\delta x=2x+x^{-2}\] \[\delta ^2 y/\delta x^2=2-2x^{-3}\]

OpenStudy (anonymous):

differentiate term by term dy/dx = 2x - (-x^-2) = 2x + x^-2 second derivative is found by a further differentiation: d2y / dx^2 = 2 - 2x^-3 or 2 - 2/x^3

OpenStudy (amistre64):

since the second derivative is just the derivative of the first derivative; id say you just take the derivative again ....